# Non-Abelian Statistics for Bosonic Symmetry-Protected Topological Phases

### v3

@@ VII Conclusions and Discussion
In summary, we discover the SPNA statistics in a bosonic SPT phase with particle-hole symmetry and particle number conservation. We show that the topological zero modes exhibit two distinct classes of braiding statistics: the conventional braiding and defect-assisted braiding, the latter being a fractionization of the former and manifesting a novel nonlinear transformation. The result can be extended to generic bosonic SPT phases with real Hamiltonian, representing a universal property of SPNA statistics across a broad class of bosonic SPT phases. A systematic theory has been developed, accompanied by a detailed study of the experimental feasibility. The predicted two classes of braiding statistics are realized through two different types of braiding schemes. For each type of braiding scheme, a concrete tri-junction has been demonstrated to braid zero modes in the bosonic SPT phase, facilitating the realization of various quantum gates. We investigate the conditions for achieving high-fidelity braiding matrices. Experimental schemes are proposed and discussed in detail for observing these predictions using analog quantum simulations.
This study unveils the universal features of SPNA statistics of topological zero modes in the bosonic SPT phases characterized by real Hamiltonians. Pertinent directions for future investigation include the SPNA statistics of topological zero modes in the more generic bosonic SPT phases, including those with complex Hamiltonians. An interesting question is whether the phase gate, which is essential for universal quantum computing, can be realized via braiding of zero modes in bosonic SPT phases. This presents an exciting avenue for the future research.
We note that this work holds the great potential to promote the first experimental realization of non-Abelian quasiparticles and their braiding statistics in quantum many-body phases. Currently, the rigorous verification of MZMs and their non-Abelian braiding in topological superconductors faces substantial challenges. Digital simulations of non-Abelian braiding are actively pursued on quantum computing platforms Google Quantum AI (2023); Xu et al. (2023); Iqbal et al. (2024), whereas realizing non-Abelian quasiparticles and their braiding statistics poses a fundamentally distinct and paramount task in topological quantum matter. The present study, with its full experimental feasibility, paves the way for the ultimate experimental observation of non-Abelian quasiparticles and their braiding statistics in the near future, which may trigger the SPT quantum computation, as a new route for topological quantum computation.
Acknowledgments
This work was supported by National Key Research and Development Program of China (2021YFA1400900), National Natural Science Foundation of China (No. 12425401 and No. 12261160368), Quantum Science and Technology-National Science and Technology Major Project (2021ZD0302000), and Shanghai Municipal Science and Technology Major Project (No. 2019SHZDZX01).
@@ Acknowledgments
APPENDIX A THE FLOQUET THEORY FOR EFFECTIVE HAMILTONIAN
@@
into Eq. (A) gives rise to the Eq. (13).
APPENDIX B BRAIDING OPERATORS FROM THE YANG-BAXTER EQUATION
In this section, we derive the braiding operators by solving the Yang-Baxter equation, thereby establishing the SPNA statistics for hard-core bosonic zero modes on a mathematically rigorous footing. We focus on the case protected by the unitary symmetries of particle-hole transformation $C$ and particle-number conservation $U_{\theta}$. For two zero modes, the most general effective Hamiltonian governing the braiding process is:
$H_{E}=\lambda G_{12}+\delta P_{12},$ (46)
where $\lambda$, $\delta\in\mathbb{R}$ ensure Hermiticity, $G_{12}\equiv\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}$ is symmetric under exchange of indices, and $P_{12}\equiv G_{12}^{2}=n_{1}+n_{2}-2n_{1}n_{2}$ is the projector onto the subspace $\{\left|{\beta_{1}}\right\rangle,\left|{\beta_{2}}\right\rangle\}$. These operators satisfy
@@
$U_{12}$ $=$ $e^{-iH_{E}T}=e^{-i\lambda TG_{12}}e^{-i\delta TP_{12}}$
$=$ $1-ie^{-i\theta}\text{sin}(\varphi)G_{12}+(e^{-i\theta}\text{cos}(\varphi)-1)P_{12},$
@@
substituting Eq. (B) into Eq. (49) gives:
$aG_{12}+2abG_{12}-aG_{23}-2abG_{23}+a^{2}P_{12}+bP_{12}+b^{2}P_{12}-a^{2}P_{23}-bP_{23}-b^{3}P_{12}P_{23}+ab^{2}P_{12}G_{23}P_{12}+a^{2}bP_{12}G_{23}G_{12}$
$-b^{2}P_{23}+a^{3}G_{12}G_{23}G_{12}+a^{2}bG_{12}G_{23}P_{12}+a^{2}bG_{12}P_{23}G_{12}+ab^{2}G_{12}P_{23}-ab^{2}P_{12}G_{23}+ab^{2}P_{23}G_{12}+b^{3}P_{23}P_{12}$
@@
where we have defined $a\equiv-ie^{-in\theta}\text{sin}(n\varphi)$ and $b\equiv e^{-in\theta}\text{cos}(\varphi)-1$, and used $P_{12},P_{23}]=0$ to simplify intermediate steps. Using the additional algebraic relations
$\{P_{12},G_{23}\}=G_{23},\{P_{23},G_{12}\}=G_{12},\{G_{12},G_{23}\}=G_{13},$
the above expression Eq. (B) reduces to
$(a+2ab$ $+$ $ab^{2})(G_{12}-G_{23})$
@@
Without loss of generality, we restrict $\theta$ and $\varphi$ to the range $(-\pi,\pi]$. The nontrivial solutions of Eq. (52) are $\varphi=\pm\frac{\pi}{2}$. The choice of the plus or minus sign is merely conventional, as the two cases are related by Hermitian conjugation accompanied by reversing the sign of $\theta$. Physically, the sign corresponds to the braiding direction—clockwise or counterclockwise. In this work, we adopt the $\varphi=-\frac{\pi}{2}$ convention, and the corresponding braiding operator is
$U_{12}=1+ie^{-i\theta}G_{12}-P_{12}.$ (53)
@@
The solution of the braiding operator in Eq. (53) to the Yang-Baxter equation depends on a continuous parameter $\theta$. This dependence originates from the term $\propto P_{12}$ in $H_{E}$, which physically describes an interaction between the two zero modes: $P_{12}=1$ when exactly one zero mode is occupied and $P_{12}=0$ otherwise. Such an interaction leads to phase accumulation during braiding, thereby producing a continuous family of solutions for the braiding operator. The reality condition of the Hamiltonian, however, imposes a further restriction, quantizing the allowed solutions and leaving only two distinct classes. In particular, a real Hamiltonian requires that the Berry phase associated with a full braiding process be either +1+1 or 1-1, namely
$U^{2}_{12}=1-(1+e^{-2i\theta})P_{12}=\pm 1.$ (54)
@@
We refer to the cases $U^{2}=1$ and $U^{2}=-1$ as conventional braiding statistics and defect-assisted braiding statistics, respectively. When both zero modes are either occupied or empty, $P_{12}=0$, yielding $U^{2}_{12}=1$. This result is physically consistent because the bosonic commutation relations of the zero-mode operators require that the singlet sectors $\left|{0,0}\right\rangle$ and $\left|{1,1}\right\rangle$ are invariant under braiding. Nevertheless, when exactly one zero mode is occupied, we have $U_{12}^{2}=-e^{-2i\theta}$. This condition leads to two distinct classes of solutions for the parameter $\theta$: (i) $\theta=\pm\pi/2$, and (ii) $\theta=0$ or $\pi$, corresponding to two classes of effective Hamiltonians. For the former, the effective Hamiltonian realizes the conventional braiding statistics, with the braiding operator
$U^{\text{conv}}_{12}=1+G_{12}-P_{12},$ (55)
@@
where we take $\theta=\pi/2$ by convention, as the opposite choice only introduces an irrelevant overall phase factor.
For the second case, we choose $\theta=0$, and the braiding operator reads
$\tilde{U}^{\text{def}}_{12}=1+iG_{12}-P_{12},$
where the opposite choice differs only by Hermitian conjugation. We note that the factor of $i$ in $\tilde{U}^{\text{def}}_{12}$ is a gauge choice, while the physical braiding operator should contain only real coefficients since the system evolves under a real Hamiltonian. This will be rigorously shown in Appendix. C. The imaginary factor can be removed by applying a gauge transformation,
$\mathcal{U}\beta_{1}\mathcal{U}^{\dagger}=i\beta_{1},\,\mathcal{U}\beta_{2}\mathcal{U}^{\dagger}=\beta_{2},$
with which we reach finally
$U^{\text{def}}_{12}\equiv\mathcal{U}^{\dagger}\tilde{U}^{\text{def}}_{12}\mathcal{U}=1+\beta_{2}^{\dagger}\beta_{1}-\beta_{1}^{\dagger}\beta_{2}-P_{12}.$ (56)
From the above analysis, we identify two key ingredients responsible for the quantization of non-Abelian statistical solutions: (i) the reality of the Hamiltonian, and (ii) the interaction between zero modes, manifested as a nonvanishing $P_{12}$ term in the effective Hamiltonian. Notably, the second ingredient arises from the strong-correlation nature of the bosonic SPT phase and does not appear naturally in free-fermion topological phases. Our results thus reflect an intrinsic feature of strongly correlated systems.
APPENDIX C THE NON-ABELIAN BERRY PHASE AND TOPOLOGICAL CLASSIFICATION
@@
We first consider defect-assisted braiding. In this case, the system is occupied by a single hard-core boson, and at $t=0$ the zero-energy subspace (the ground-state subspace in the presence of a local defect) is spanned by the basis states $(\left|{1}\right\rangle,\left|{4}\right\rangle)$, where $\left|{j}\right\rangle\equiv b_{j}^{\dagger}\left|{\Uparrow}\right\rangle$. To ensure smooth evolution of the basis vectors, we adopt the following gauge choice:
@@
APPENDIX D CALCULATIONS IN THE DIMERIZED LIMIT
Figure 8: Numerical time evolution of zero-mode wavefunctions for two braiding schemes in the dimerized limit. (a) Amplitudes of and (b) phase difference between $\left|{\beta_{1}(t)}\right\rangle$ and $\left|{\beta_{2}(t)}\right\rangle$ for the first type of braiding scheme. (c) Amplitudes and (d) phase difference for the second type of braiding scheme. The horizontal axis represents the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0$. Simulations are performed with a time step $\Delta t=0.1$ and an elementary step time $T_{0}=80$. Each braiding involves $9$ elementary steps of zero-mode motion, yielding a total braiding time $T=9T_{0}=720$. The red horizontal line indicates the theoretical prediction of $\Delta\phi=\text{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\text{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\text{mod}\,2\pi$ for both schemes. The vertical dashed line marks the completion of the braiding process, after which the system is held static to show the stabilized value of $\Delta\phi$. The numerical results of $\Delta\phi$ for both classes of statistics are indicated in the figure.
We first compute the geometric phase of an elementary moving step illustrated in Fig. 1(e). The dynamics of an elementary moving step is governed by the Hamiltonian $H_{\text{e}}$ in Eq. (23). In the dimerized limit, the parameters at $t=0$ are initialized with $u(0)=0$, $v(0)=v$. Adiabatically tuning $u(t)$ to $u(T_{0})=v$ and $v(t)$ to $v(T_{0})=0$ moves the zero mode from site $1$ to site $3$. The geometric phase obtained by the zero mode depends on the particle-number occupation on sites $2$ and $3$. In both the single-particle subspace spanned by $\{b_{1}^{\dagger}|\Uparrow\rangle,b_{2}^{\dagger}|\Uparrow\rangle,b_{3}^{\dagger}|\Uparrow\rangle\}$ and the two-particle subspace spanned by $\{b_{2}^{\dagger}b_{3}^{\dagger}|\Uparrow\rangle,b_{3}^{\dagger}b_{1}^{\dagger}|\Uparrow\rangle,b_{1}^{\dagger}b_{2}^{\dagger}|\Uparrow\rangle\}$, the $H_{\text{e}}$ is
$H_{\text{e}}=\begin{pmatrix}0&u(t)&0\\ u(t)&0&v(t)\\ 0&v(t)&0\end{pmatrix}$. (71)
$b_{1}^{\dagger}|\Uparrow\rangle$ $\to$ $-b_{3}^{\dagger}|\Uparrow\rangle,$ (72)
$b_{1}^{\dagger}(b_{2}^{\dagger}-b_{3}^{\dagger})|\Uparrow\rangle$ $\to$ $b_{3}^{\dagger}(b_{2}^{\dagger}-b_{1}^{\dagger})|\Uparrow\rangle.$ (73)
$H(t)=\sum_{i=1,2,...,6,8,9}v_{i}(t)b^{\dagger}_{i+1}b_{i}+v_{7}(t)b^{\dagger}_{8}b_{4}+h.c..$ (74)
In the braiding we adjust the coupling strengths over time following the functions depicted in Fig. 7(c). The rising (falling) slope in Fig. 7(c) follow the form $\chi(t/T_{0})$ ($1-\chi(t/T_{0})$) defined in Eq. (25). The numerical results, shown in Fig. 8, match the theoretical predictions.
APPENDIX E ERROR PREVENTION THROUGH LOCAL DEFECT ISOLATION
Figure 9: Imperfect defect-assisted braiding procedure. The golden dashed lines represent the residual coupling strength $\delta$ between the zero modes and other parts of the tri-junction as the zero modes pass through the local defect.
Defect-assisted braiding requires isolating the local defect to ensure that it returns to the original state after braiding. Failure to meet this condition causes errors in the braiding results. In the following, we analyze these errors and verify that our scheme through isolating the local defect effectively prevents them. Specifically, we examine an imperfect defect-assisted braiding procedure where the zero mode is not fully localized before and after passing through the local defect, but retains a residual coupling $\delta$ with a portion of the tri-junction, as shown in Fig. 9. We numerically compute the change in average fidelity as a function of the residual coupling strength $\delta$. The results, shown in Fig. 10, confirm that our scheme (with $\delta=0$) effectively prevents errors.
We also calculate the unitarity (defined in Appendix F) of the braiding matrix. As shown in Fig. 10, the unitarity follows a behavior similar to that of the average fidelity, confirming that errors arise when the local defect fails to fully return to its original state. This leads to components outside the subspace $\mathcal{H}_{0}$ in the states $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$, resulting in a non-unitary braiding matrix. Setting $\delta=0$ in perfect defect-assisted braiding solves this problem. It also suggests that measuring the local defect state offers a practical method to assess braiding results.
APPENDIX F AVERAGE FIDELITY AND UNITARITY
Figure 10: Numerical calculations of the average fidelity $F$ and unitarity $\tilde{U}$ as functions of the residual coupling strength $\delta$ (in the unit of maximal coupling strength $v$) for imperfect defect-assisted braiding procedures.

### v4

@@ VII Conclusions and Discussion
In summary, we discover a new paradigm of SPNA statistics in a representative bosonic SPT phase protected by unitary symmetry, with a systematic and rigorous theory being established. The topological zero modes exhibit two distinct classes of braiding statistics: the conventional and defect-assisted unconventional braiding statistics, the latter being a fractionization of the former and manifesting a novel nonlinear transformation. The braiding statistics can be extended to generic bosonic SPT phases with real Hamiltonian, representing a fundamental dynamical character of such topological phases. We propose feasible tri-junctions with or without the assistance of a controlled defect to realize the two classes of braiding statistics in the real systems, with which the non-commutative quantum gates are achieved in the single-qubit setting formed by two zero modes. The results also facilitate the realization of various single- and two-qubit quantum gates. Experimental schemes are proposed and discussed in detail for observation of these predictions with high-fidelity, paving the way for the experimental study in the near future.
Pertinent directions for future investigation include the SPNA statistics of topological zero modes in the more generic bosonic SPT phases, including those with complex Hamiltonians. An interesting question is whether the phase gate, which is essential for universal quantum computing, can be realized via braiding of zero modes in bosonic SPT phases. This presents an exciting avenue for the future research.
We note that this work holds the great potential to promote the first experimental realization of non-Abelian quasiparticles and their braiding statistics in quantum many-body phases. Currently, the rigorous verification of MZMs and their non-Abelian braiding in topological superconductors faces substantial challenges. Digital simulations of non-Abelian braiding are actively pursued on quantum computing platforms Google Quantum AI (2023); Xu et al. (2023); Iqbal et al. (2024), whereas realizing non-Abelian quasiparticles and their braiding statistics poses a fundamentally distinct and paramount task in topological quantum matter. The present study, with its full experimental feasibility, paves the way for the ultimate experimental observation of non-Abelian quasiparticles in topological quantum phases and their braiding statistics in the near future, which may trigger the SPT quantum computation, as a new route for topological quantum computation.
@@
VIII Methods
VIII.1 The floquet theory for effective Hamiltonian
@@
into Eq. (VIII.1) gives rise to the Eq. (11).
VIII.2 Braiding operators from the Yang-Baxter equation
In this section, we derive the braiding operators by solving the Yang-Baxter equation, thereby establishing the SPNA statistics for hard-core bosonic zero modes on a mathematically rigorous footing. We focus on the case protected by the unitary symmetries of particle-hole transformation and particle-number conservation. For two zero modes, the most general effective Hamiltonian governing the braiding process is:
$H_{E}=\lambda G_{12}(\alpha)+\delta P_{12},$ (49)
where $\lambda$, $\delta\in\mathbb{R}$, $G_{12}(\alpha)\equiv e^{i\alpha}\beta_{1}^{\dagger}\beta_{2}+e^{-i\alpha}\beta_{2}^{\dagger}\beta_{1}$ with $\alpha\in[0,2\pi)$, and $P_{12}\equiv G_{12}^{2}=n_{1}+n_{2}-2n_{1}n_{2}$ is the projector onto the half-filled subspace $\mathcal{H}_{0}$. These operators satisfy
@@ VIII.2 Braiding operators from the Yang-Baxter equation
$U_{12}$ $=$ $e^{-iH_{E}T}=e^{-i\lambda TG_{12}(\alpha)}e^{-i\delta TP_{12}}$
$=$ $I-ie^{-i\theta}\text{sin}(\varphi)G_{12}(\alpha)+(e^{-i\theta}\text{cos}(\varphi)-1)P_{12},$
@@
substituting Eq. (VIII.2) into Eq. (52) gives:
$aG_{12}+2abG_{12}-aG_{23}-2abG_{23}+a^{2}P_{12}+bP_{12}+b^{2}P_{12}-a^{2}P_{23}-bP_{23}+ab^{2}P_{12}G_{23}P_{12}+a^{2}bP_{12}G_{23}G_{12}$
$-b^{2}P_{23}+a^{3}G_{12}G_{23}G_{12}+a^{2}bG_{12}G_{23}P_{12}+a^{2}bG_{12}P_{23}G_{12}+ab^{2}G_{12}P_{23}-ab^{2}P_{12}G_{23}+ab^{2}P_{23}G_{12}$
@@
where we have defined $a\equiv-ie^{-i\theta}\text{sin}(\varphi)$ and $b\equiv e^{-i\theta}\text{cos}(\varphi)-1$, and used $P_{12},P_{23}]=0$ to simplify intermediate steps. Using the additional algebraic relations
$\{P_{12},G_{23}(\alpha)\}$ $=$ $G_{23}(\alpha),$ (54)
$\{P_{23},G_{12}(\alpha)\}$ $=$ $G_{12}(\alpha),$ (55)
the above expression Eq. (VIII.2) reduces to
$(a+2ab$ $+$ $ab^{2})(G_{12}(\alpha)-G_{23}(\alpha))$
@@
For simplicity and without loss of physical content, we restrict $\theta$ and $\varphi$ to the range $(-\pi,\pi]$; more general solutions are $\theta+2n\pi$ and $\varphi+2m\pi$ with $m,n\in\mathbb{Z}$. The nontrivial solutions of Eq. (57) are $\varphi=\pm\frac{\pi}{2}$. The choice of the plus or minus sign is merely a convention, as the two cases are related by Hermitian conjugation accompanied by reversing the sign of $\theta$. Physically, the sign corresponds to the braiding direction—clockwise or counterclockwise. In this work, we adopt the $\varphi=-\frac{\pi}{2}$ convention, and the corresponding braiding operator is
$U_{12}=I+ie^{-i\theta}G_{12}(\alpha)-P_{12}.$ (58)
@@
The solution of the braiding operator in Eq. (58) to the Yang-Baxter equation depends on a continuous parameter $\theta$. This dependence originates from the term $\propto P_{12}$ in $H_{E}$, which physically describes an interaction between the two zero modes: $P_{12}=1$ when exactly one zero mode is occupied and $P_{12}=0$ otherwise. Such an interaction leads to phase accumulation during braiding, thereby producing a continuous family of solutions for the braiding operator. The fact that the Hamiltonian is real imposes an additional constraint: the Berry phase must be real. Consequently, for a full braiding process, the corresponding transformation is
$U^{2}_{12}=1-(1+e^{-2i\theta})P_{12}=\pm I.$ (59)
@@
We refer to the cases $U^{2}=I$ and $U^{2}=-I$ as conventional braiding statistics and defect-assisted braiding statistics, respectively. When both zero modes are either occupied or empty, $P_{12}=0$, yielding $U^{2}_{12}=1$. This result is physically consistent because the bosonic commutation relations of the zero-mode operators require that the singlet sectors $\left|{0,0}\right\rangle$ and $\left|{1,1}\right\rangle$ are invariant under braiding. Nevertheless, when exactly one zero mode is occupied, we have $U_{12}^{2}=-e^{-2i\theta}$. This condition leads to two distinct classes of solutions for the parameter $\theta$: (i) $\theta=\pm\pi/2$, and (ii) $\theta=0$ or $\pi$, corresponding to two classes of effective Hamiltonians. Moreover, as shown in Fig. 3 and in Methods VIII.3, the braiding transformation of the ground states is real. This requires that the coefficients in Eq. (58) be real, which in turn fixes $\alpha$ with a given $\theta$. For the firt class, the effective Hamiltonian realizes the conventional braiding statistics, with the braiding operator
$U^{\text{conv}}_{12}$ $=$ $I+G_{12}(0)-P_{12},$ (60)
$=$ $I+\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2},$
@@
where we take $\theta=\pi/2$ by convention, as the opposite choice only introduces an overall 1-1 factor to the braiding matrix in the subspace $\mathcal{H_{0}}$. Since the coefficients in Eq. (60) are already real, the $\alpha$ takes the value 0 (Also, $\alpha=\pi$ only contributes an overall factor of 1-1).
For the second class, we choose $\theta=0$ with the associated $\alpha=\frac{\pi}{2}$ to ensure that the braiding operator is real. It then reads
$U^{\text{def}}_{12}$ $=$ $I+iG_{12}(\frac{\pi}{2})-P_{12},$ (61)
$=$ $I-\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2},$
where the opposite choice differs only by Hermitian conjugation. These results, Eqs. (60) and (61), are confirmed with the numerical results and the exact solution of the wavefunction evolution during the braiding process, in which all the coefficients are real, as shown in the following Methods VIII.3.
VIII.3 The non-Abelian Berry phase and topological classification
@@
We first consider defect-assisted braiding. In this case, the system is occupied by a single hard-core boson, and at $t=0$ the zero-energy subspace (the ground-state subspace in the presence of a local defect) is spanned by the basis states $(\left|{1}\right\rangle,\left|{4}\right\rangle)$, where $\left|{j}\right\rangle\equiv b_{j}^{\dagger}\left|{\Uparrow}\right\rangle$ with $\left|{\Uparrow}\right\rangle$ denoting the state in which all spins are in the up state. To ensure smooth evolution of the basis vectors, we adopt the following gauge choice:
@@
VIII.4 Average fidelity and unitarity
@@
Data availability.–The data that support the findings of this study are available from the corresponding author upon reasonable request.
Code availability.–All numerical codes in this paper are available upon request to the authors.
Supplementary Material
S1 CALCULATIONS IN THE DIMERIZED LIMIT
We first compute the geometric phase of an elementary moving step illustrated in Fig. 1(e). The dynamics of an elementary moving step is governed by the Hamiltonian $H_{\text{e}}$ in Eq. (24). In the dimerized limit, the parameters at $t=0$ are initialized with $u(0)=0$, $v(0)=v$. Adiabatically tuning $u(t)$ to $u(T_{0})=v$ and $v(t)$ to $v(T_{0})=0$ moves the zero mode from site $1$ to site $3$. The geometric phase obtained by the zero mode depends on the particle-number occupation on sites $2$ and $3$. In both the single-particle subspace spanned by $\{b_{1}^{\dagger}|\Uparrow\rangle,b_{2}^{\dagger}|\Uparrow\rangle,b_{3}^{\dagger}|\Uparrow\rangle\}$ and the two-particle subspace spanned by $\{b_{2}^{\dagger}b_{3}^{\dagger}|\Uparrow\rangle,b_{3}^{\dagger}b_{1}^{\dagger}|\Uparrow\rangle,b_{1}^{\dagger}b_{2}^{\dagger}|\Uparrow\rangle\}$, where $\left|{\Uparrow}\right\rangle$ denotes the state with all spins up, the $H_{\text{e}}$ is
$H_{\text{e}}=\begin{pmatrix}0&u(t)&0\\ u(t)&0&v(t)\\ 0&v(t)&0\end{pmatrix}$. (S1)
$b_{1}^{\dagger}|\Uparrow\rangle$ $\to$ $-b_{3}^{\dagger}|\Uparrow\rangle,$ (S2)
$b_{1}^{\dagger}(b_{2}^{\dagger}-b_{3}^{\dagger})|\Uparrow\rangle$ $\to$ $b_{3}^{\dagger}(b_{2}^{\dagger}-b_{1}^{\dagger})|\Uparrow\rangle.$ (S3)
$H(t)=\sum_{i=1,2,...,6,8,9}v_{i}(t)b^{\dagger}_{i+1}b_{i}+v_{7}(t)b^{\dagger}_{8}b_{4}+h.c..$ (S4)
In the braiding we adjust the coupling strengths over time following the functions depicted in Fig. 7(c). The rising (falling) slope in Fig. 7(c) follow the form $\chi(t/T_{0})$ ($1-\chi(t/T_{0})$) defined in Eq. (26). The numerical results, shown in Fig. S1, match the theoretical predictions.
Figure S1: Numerical time evolution of zero-mode wavefunctions for two braiding schemes in the dimerized limit. (a) Amplitudes of and (b) phase difference between $\left|{\beta_{1}(t)}\right\rangle$ and $\left|{\beta_{2}(t)}\right\rangle$ for the first type of braiding scheme (conventional braiding). (c) Amplitudes and (d) phase difference for the second type of braiding scheme (defect-assisted braiding). The horizontal axis represents the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0$. Simulations are performed with a time step $\Delta t=0.1$ and an elementary step time $T_{0}=80$. Each braiding involves $9$ elementary steps of zero-mode motion, yielding a total braiding time $T=9T_{0}=720$. The red horizontal line indicates the theoretical prediction of $\Delta\phi=\text{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\text{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\text{mod}\,2\pi$ for both schemes. The vertical dashed line marks the completion of the braiding process, after which the system is held static to show the stabilized value of $\Delta\phi$. The numerical results of $\Delta\phi$ for both classes of statistics are indicated in the figure.
S2 ERROR PREVENTION THROUGH LOCAL DEFECT ISOLATION
Figure S2: Imperfect defect-assisted braiding procedure. The golden dashed lines represent the residual coupling strength $\delta$ between the zero modes and other parts of the tri-junction as the zero modes pass through the local defect.
Defect-assisted braiding requires isolating the local defect to ensure that it returns to the original state after braiding. Failure to meet this condition causes errors in the braiding results. In the following, we analyze these errors and verify that our scheme through isolating the local defect effectively prevents them. Specifically, we examine an imperfect defect-assisted braiding procedure where the zero mode is not fully localized before and after passing through the local defect, but retains a residual coupling $\delta$ with a portion of the tri-junction, as shown in Fig. S2. We numerically compute the change in average fidelity as a function of the residual coupling strength $\delta$. The results, shown in Fig. S3, confirm that our scheme (with $\delta=0$) effectively prevents errors.
We also calculate the unitarity (defined in Methods VIII.4) of the braiding matrix. As shown in Fig. S3, the unitarity follows a behavior similar to that of the average fidelity, confirming that errors arise when the local defect fails to fully return to its original state. This leads to components outside the subspace $\mathcal{H}_{0}$ in the states $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$, resulting in a non-unitary braiding matrix. Setting $\delta=0$ in perfect defect-assisted braiding solves this problem. It also suggests that measuring the local defect state offers a practical method to assess braiding results.
Figure S3: Numerical calculations of the average fidelity $F$ and unitarity $\tilde{U}$ as functions of the residual coupling strength $\delta$ (in the unit of maximal coupling strength $v$) for imperfect defect-assisted braiding procedures.